# Efficient and Accurate Optimal Transport with Mirror Descent and Conjugate Gradients

### v3

@@ 4.3 Preconditioned Non-linear Conjugate Gradients for Bregman Projections
SK converges more slowly at low temperatures (Kosowsky & Yuille, 1994). [^3] In Appx. A we provide a proposition that applies Thm. 4 of Franklin & Lorenz (1989) on matrix scaling to the MD setting, which also characterizes this behavior, albeit asymptotically. For a faster alternative, we develop Alg. 2 based on non-linear CG (NCG) methods (Fletcher & Reeves, 1964; Nocedal & Wright, 2006), which we now briefly review. Given an objective $g$, NCG takes descent directions ${{\bm{p}}^{(0)}=-\nabla g({\bm{z}}^{(0)})}$ and ${{\bm{p}}^{(k)}\leftarrow-\nabla g^{(k)}+\beta^{(k)}{\bm{p}}^{(k-1)}}$, and iterates ${{\bm{z}}^{(k+1)}\leftarrow{\bm{z}}^{(k)}+\alpha^{(k)}{\bm{p}}^{(k)}}$, where $\alpha^{(k)}$ is the step size. Optimal $\alpha^{(k)}$ has a closed-form solution for quadratics, but for general non-linear objectives, line search is necessary to find suitable step sizes $\alpha^{(k)}$. Various formulas for computing $\beta^{(k)}$ exist; for quadratic objectives, they are equivalent and guarantee convergence in at most $n^{\prime}$ iterations, where ${n^{\prime}\leq n}$ is the number of distinct eigenvalues of $\nabla^{2}g$. Further, the objective decreases faster if eigenvalues are tightly clustered (Stiefel, 1958; Kaniel, 1966; Nocedal & Wright, 2006). For example, the Hestenes-Stiefel formula sets (Nocedal & Wright, 2006):
$\tilde{{\bm{s}}}=-\mathbf{D}\big{(}\mathbf{diag}(\nabla^{2}g)\big{)}^{-1}\nabla g=\left(\frac{{\bm{r}}-{\bm{r}}(P)}{{\bm{r}}(P)},\frac{{\bm{c}}-{\bm{c}}(P)}{{\bm{c}}(P)}\right)=\left(\frac{{\bm{r}}}{{\bm{r}}(P)},\frac{{\bm{c}}}{{\bm{c}}(P)}\right)-\bm{1}_{2n}.$ (15)
@@
The Sinkhorn direction can be understood as the result of an alternative diagonal preconditioner, namely $M=\mathbf{D}\left(-\nabla g/{\bm{s}}\right)$, since ${\bm{s}}=-M^{-1}\nabla g$. Furthermore, for any sub-optimal $({\bm{u}},{\bm{v}})$, we have
@@
Figure 3: Comparison of BregmanProject algorithms (used in L7 of Alg. 1) over the upsampled MNIST dataset with $L_{1}$ (left) and $L_{2}^{2}$ (right) costs. Algorithms evaluated are the SK algorithm, the newly proposed PNCG given in Alg. 2, and a variant (see text). In all 36 problems, ${n=4096,p=1.5}$ and ${q=2}$. Each curve shows the convergence behavior, at one specific temperature $1/\gamma$, in terms of the median number of LogSumExp reductions (x-axis) until gradient norm (y-axis) reaches below target dual gradient norm $\varepsilon_{\mathrm{d}}(\gamma)$.
@@ 5 EXPERIMENTS
In this section, we first detail the MNIST experimental setup in Figs. 1-3. Then, we describe an additional color transfer task we use for benchmarking. Next, performance evaluations in terms of precision vs. wall-clock time are discussed given the results over 4 sets of problems shown in Fig. 4. In Appx. F, we add 20 more problem sets from the DOTmark benchmark of Schrieber et al. (2017) showing similar results both in terms of wall-clock time and operation counts. Lastly, the dependence on problem size $n$ is investigated in Fig. 5. All experiments were performed on an NVIDIA GeForce RTX 2080 Ti GPU with 64-bit precision.
@@ 5.1 Experimental Setup
Figure 4: Wall-clock time vs. error benchmarking over the upsampled MNIST (top) and color transfer (bottom) problems using $L_{1}$ (left) and $L_{2}^{2}$ (right) distances as cost functions. Each marker shows the median time to converge (over 18 random problems) for each algorithm at a given hyperparameter setting, which controls the precision level, and the error $\langle P-P^{*},C\rangle$ after rounding the output of the algorithm onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$ – with the exception of Alg. 3.5 of Feydy (2020); see text.
@@ 5.2 Wall-clock Time Comparisons With Prior Work
In Fig. 4, we present wall-clock time benchmarking of MDOT (with both Sinkhorn and PNCG projections) against existing algorithms on the upsampled MNIST and color transfer problems. All benchmark methods were implemented in PyTorch and run on the GPU. For MDOT, we use $q{=}2^{1/3},p=1.5$ and $\gamma_{\mathrm{i}}{=}2^{4}$ in all experiments. For the closely related Mirror Sinkhorn (MSK) algorithm of Ballu & Berthet (2023) the variable step size schedule prescribed by their Thm. 3.3 is used in our implementation. For Alg. 3.5 of Feydy (2020), we decay temperature at a rate $q=0.7^{-1}$, which interpolates their fast ($q{=}0.5^{-1}$) and safe ($q{=}0.9^{-1}$) settings. For AAM (Guminov et al., 2021), Mirror Prox Sherman Optimized (Jambulapati et al., 2019) and APDAGD (Dvurechensky et al., 2018), each implementation closely follows an open-source NumPy implementation. Our PyTorch implementation was verified to produce identical results to the publicly available NumPy code. We additionally attempted comparison with APDAMD (Lin et al., 2019) and PDASMD (Luo et al., 2023), but observed extremely long convergence times for $n=4096$ and omitted the results. For further details on the implementation of benchmark methods, we refer the reader to Appx. E.
@@
While MDOT optimizes (2.1) to satisfy a convergence criterion following each temperature decrease, Alg. 3.5 of Feydy (2020) performs a single (symmetrized) Sinkhorn update instead, i.e., it does not minimize the sequence of dual objectives sufficiently despite taking increasingly large gradient steps in the dual space (cf. 4-5). This causes an accumulation of projection errors and results in the algorithm hitting a precision wall. Their debiasing option for estimating the OT distance via Sinkhorn divergences (introduced by Ramdas et al. (2017)) fares slightly better and is used here to comprise a stronger baseline, albeit this approach does not find a member of ${\mathcal{U}}({\bm{r}},{\bm{c}})$, which may be a strict requirement in some applications. MSK also runs a single row/column scaling update after a temperature decrease, but takes increasingly smaller steps and maintains a running average of transport plans to ensure convergence. It performs well at low precision, but shrinking step sizes slow it down, so that it exhibits $O(n^{2}\varepsilon^{-2})$ convergence behavior. Sinkhorn iteration (log-domain stabilized, see Alg. 4 in the Appx. A) benefits substantially from setting $p=1.5$ rather than $p=1$ at sufficiently low temperatures for $L_{1}$ costs (see also Sec. 4.2.2). APDAGD underperforms SK with ${p=1}$ and AAM performs similarly to it. Mirror Prox Sherman Optimized of Jambulapati et al. (2019) overtakes SK (${p=1.0}$) in one case only (top-left) in the high precision range. Meanwhile, MDOT-Sinkhorn enjoys faster convergence than the more competitive SK (${p=1.5}$) owing to warm-started temperature annealing, especially in the high precision range (near the solution). MDOT-PNCG is the quickest to converge in all cases. The performance gap with its close second, MDOT-Sinkhorn, grows with higher precision.
@@ 5.3 Empirical Dependence on Problem Size of MDOT-PNCG
Our last set of experiments investigates the practical dependence of MDOT-PNCG on the problem size $n$. Over the same 4 problem sets as Fig. 4, we change $n$ from $36$ to $16,384$ by up- or down-sampling images. The $n$ values are selected to be approximately equally spaced on a logarithmic scale. In Fig. 5, we plot the behavior of MDOT-PNCG for a range of final temperature values ${\gamma_{\mathrm{f}}\in\{2^{6},2^{9},2^{12},2^{15}\}}$. At medium precision (green and blue), we find that the algorithm behaves no worse than $O(n^{2})$ in practice as implied by the flatness of the curves. As seen visually, at higher precision (roughly 5-decimals) with ${\gamma_{\mathrm{f}}\in\{2^{12},2^{15}\}}$, the proposed GPU-parallel algorithm behaves roughly as $O(n^{5/2})$ at worst and even better for some of the problems in practice. These should be compared to the $\widetilde{O}(n^{5/2})$ theoretical and $\widetilde{O}(n^{3})$ practical complexity of CPU-based exact solvers (Pele & Werman, 2009; Lee & Sidford, 2014).
Figure 5: Problem size dependence of MDOT-PNCG convergence over the upsampled MNIST (top) and color transfer (bottom) problems using $L_{1}$ (left) and $L_{2}^{2}$ (right) distance cost functions with ${\gamma_{\mathrm{f}}\in\{2^{6},2^{9},2^{12},2^{15}\}}$. Each marker displays the median over 20 problems and shaded areas show 75% confidence intervals.
@@ 6 CONCLUSION
In this work, we first presented a general procedure, MDOT, for computing OT distances with high precision and described its relation to a well-known temperature annealing strategy ($\varepsilon$-scaling). MDOT employs a novel warm-starting of the sequence of EOT dual problems encountered in temperature annealing, which was empirically shown to be highly effective compared to existing approaches. In addition, a specialized non-linear CG algorithm was developed as an alternative to Sinkhorn iteration and was shown to be more effective at low temperatures (under weak regularization). Over 24 different problem sets, the combined MDOT-PNCG algorithm outperforms aggressively tuned Sinkhorn iteration and many other recent baselines in terms of convergence of the primal optimality gap measured in wall-clock time. The algorithm was also shown to behave well with respect to the problem size. Interesting directions for future research include the theoretical convergence behavior of PNCG and bounds on the gradient norm of our warm-started initialization, as well as the development of faster Bregman projection algorithms and warm-starting methods. Adaptive temperature decay and dual problem stopping criteria are also of interest.
@@
Acknowledgements. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
A Proofs
A.1 Derivation of dual objective minimization as Bregman projection
For the interested reader, we provide here a short sketch of the relationship
$P({\bm{u}}^{*},{\bm{v}}^{*};\gamma)=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\big{(}P|P({\bm{u}}^{\prime},{\bm{v}}^{\prime};\gamma)\big{)},$ (19)
given any initial ${\bm{u}}^{\prime},{\bm{v}}^{\prime}\in\mathbb{R}^{n}$ and ${\bm{u}}^{*},{\bm{v}}^{*}\in\operatorname*{arg\,min}_{{\bm{u}},{\bm{v}}\in\mathbb{R}^{n}}g({\bm{u}},{\bm{v}};\gamma)$. First, given the definition in Sec. 2 of $D_{\mathrm{KL}}$ for un-normalized $P\in\mathbb{R}^{n\times n}_{>0}$, observe that

### v4

@@ 4.3 Preconditioned Non-linear Conjugate Gradients for KL Projections
This has the benefit that it can be evaluated via numerically stable LogSumExp reductions, e.g., see lines 2-3 of Alg. 2. The Sinkhorn direction can be understood as the result of an alternative diagonal preconditioner, namely $M=\mathbf{D}\left(-\nabla g/{\bm{s}}\right)$, since ${\bm{s}}=-M^{-1}\nabla g$. Furthermore, for any sub-optimal $({\bm{u}},{\bm{v}})$, we have
@@
Figure 4: Comparison of KL projection algorithms (used in L7 of Alg. 1) over the upsampled MNIST dataset with $L_{1}$ (left) and $L_{2}^{2}$ (right) costs. Algorithms evaluated are the SK algorithm, the newly proposed PNCG given in Alg. 2, and a variant (see text). In all 36 problems, ${n=4096,p=1.5}$ and ${q=2}$. Each curve shows the convergence behavior, at one specific temperature $1/\gamma$, in terms of the median number of LogSumExp reductions (x-axis) until gradient norm (y-axis) reaches below target dual gradient norm $\varepsilon_{\mathrm{d}}(\gamma)$.
@@ 5 EXPERIMENTS
In this section, we first detail the MNIST experimental setup in Figs. 2-4. Then, we describe an additional color transfer task we use for benchmarking. Next, performance evaluations in terms of precision vs. wall-clock time are discussed given the results over 4 sets of problems shown in Fig. 5. In Appx. H, we add 20 more problem sets from the DOTmark benchmark of Schrieber et al. (2017) showing similar results both in terms of wall-clock time and operation counts. Lastly, the dependence on problem size $n$ is investigated in Fig. 6. All experiments were performed on an NVIDIA GeForce RTX 2080 Ti GPU with 64-bit precision. Exact costs $\langle P^{*},C\rangle$ are evaluated using the implementation of the CPU-based algorithm of Bonneel et al. (2011) from the Python Optimal Transport (POT) library (Flamary et al., 2021), which was run on an Intel Xeon Silver 4110 (2.10GHz) CPU.
@@ 5.1 Experimental Setup
Figure 5: Wall-clock time vs. error benchmarking over the upsampled MNIST (top) and color transfer (bottom) problems using $L_{1}$ (left) and $L_{2}^{2}$ (right) distances as cost functions. Each marker shows the median time to converge (over 18 random problems) for each algorithm at a given hyperparameter setting, which controls the precision level, and the error $\langle P-P^{*},C\rangle$ after rounding the output of the algorithm onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$ – with the exception of Alg. 3.5 of Feydy (2020); see text. Vertical dashed lines show the median time taken by the CPU-based exact solver with $80\%$ confidence intervals.
@@ 5.2 Wall-clock Time Comparisons With Prior Work
In Fig. 5, we present wall-clock time benchmarking of MDOT (with both Sinkhorn and PNCG projections) against existing GPU-parallel algorithms on the upsampled MNIST and color transfer problems. All benchmark methods were implemented in PyTorch and run on the GPU. For MDOT, we use $q{=}2^{1/3},p=1.5$ and $\gamma_{\mathrm{i}}{=}2^{4}$ in all experiments. For the closely related Mirror Sinkhorn (MSK) algorithm of Ballu & Berthet (2023) the variable step size schedule prescribed by their Thm. 3.3 is used in our implementation. For Alg. 3.5 of Feydy (2020), we decay temperature at a rate $q=0.7^{-1}$, which interpolates their fast ($q{=}0.5^{-1}$) and safe ($q{=}0.9^{-1}$) settings. For AAM (Guminov et al., 2021), Mirror Prox Sherman Optimized (Jambulapati et al., 2019) and APDAGD (Dvurechensky et al., 2018), each implementation closely follows an open-source NumPy implementation. Our PyTorch implementation was verified to produce identical results to the publicly available NumPy code. We additionally attempted comparison with APDAMD (Lin et al., 2019) and PDASMD (Luo et al., 2023), but observed extremely long convergence times for $n=4096$ and omitted the results. For further details on the implementation of benchmark methods, see Appx. F.
@@
While MDOT optimizes (2.1) to satisfy a convergence criterion following each temperature decrease, Alg. 3.5 of Feydy (2020) performs a single (symmetrized) Sinkhorn update instead, i.e., it does not minimize the sequence of dual objectives sufficiently despite taking increasingly large gradient steps in the dual space (cf. 3a-3b). This causes an accumulation of projection errors and results in the algorithm hitting a precision wall. Their debiasing option for estimating the OT distance via Sinkhorn divergences (introduced by Ramdas et al. (2017)) fares slightly better and is used here to comprise a stronger baseline, albeit this approach does not find a member of ${\mathcal{U}}({\bm{r}},{\bm{c}})$, which may be a strict requirement in some applications. MSK also runs a single row/column scaling update after a temperature decrease, but takes increasingly smaller steps and maintains a running average of transport plans to ensure convergence. It performs well at low precision, but shrinking step sizes slow it down, so that it exhibits $O(n^{2}\varepsilon^{-2})$ convergence behavior. Sinkhorn iteration (log-domain stabilized, see Alg. 4 in the Appx. A) benefits substantially from setting $p=1.5$ rather than $p=1$ at sufficiently low temperatures for $L_{1}$ costs (see also Sec. 4.2.2). APDAGD underperforms SK with ${p=1}$ and AAM performs similarly to it. Mirror Prox Sherman Optimized of Jambulapati et al. (2019) overtakes SK (${p=1.0}$) in one case only (top-left) in the high precision range. Meanwhile, MDOT-Sinkhorn enjoys faster convergence than the more competitive SK (${p=1.5}$) owing to warm-started temperature annealing, especially in the high precision range (near the solution). MDOT-PNCG is the quickest to converge in all cases. The performance gap with its close second, MDOT-Sinkhorn, grows with higher precision.
@@ 5.3 Empirical Dependence on Problem Size of MDOT-PNCG
Our last set of experiments investigates the practical dependence of MDOT-PNCG on the problem size $n$. Over the same 4 problem sets as Fig. 5, we change $n$ from $36$ to $16,384$ by up- or down-sampling images. The $n$ values are selected to be approximately equally spaced on a logarithmic scale. In Fig. 6, we plot the behavior of MDOT-PNCG for a range of final temperature values ${\gamma_{\mathrm{f}}\in\{2^{6},2^{9},2^{12},2^{15}\}}$. At medium precision (green and blue), we find that the algorithm behaves no worse than $O(n^{2})$ in practice as implied by the flatness of the curves. As seen visually, with higher precision (roughly 5-decimals) with ${\gamma_{\mathrm{f}}\in\{2^{12},2^{15}\}}$, the proposed GPU-parallel algorithm behaves roughly as $O(n^{5/2})$ at worst and even better for some of the problems in practice (see the reference line in Fig. 6). These should be compared to the $\widetilde{O}(n^{5/2})$ theoretical and $\widetilde{O}(n^{3})$ practical complexity of CPU-based exact solvers (Pele & Werman, 2009; Lee & Sidford, 2014). Indeed, Figure 5 suggests that the CPU-based solver offers a strictly better precision-time trade-off beyond an optimality gap of around $10^{-4}{-}10^{-5}$ for this value of $n=4096$ and this particular CPU-GPU setup; however, in Table 1 of the Appendix, we show that the trade-off skews in favor of MDOT-PNCG when $n$ is increased.
Figure 6: Problem size dependence of MDOT-PNCG convergence over the upsampled MNIST (top) and color transfer (bottom) problems using $L_{1}$ (left) and $L_{2}^{2}$ (right) distance cost functions with ${\gamma_{\mathrm{f}}\in\{2^{6},2^{9},2^{12},2^{15}\}}$. Each marker displays the median over 20 problems and shaded areas show 75% confidence intervals. Dashed lines show $f(n)=an^{5/2}$ for visual comparison, with $a$ selected to intersect the purple curve at the lowest $n$.
@@ 6 CONCLUSION
In this work, we first presented a general procedure, MDOT, for computing OT distances with high precision and described its relation to a well-known temperature annealing strategy ($\varepsilon$-scaling). MDOT employs a novel warm-starting of the sequence of EOT dual problems encountered in temperature annealing, which was empirically shown to be highly effective compared to existing approaches. In addition, a specialized non-linear CG algorithm was developed as an alternative to Sinkhorn iteration and was shown to be more effective at low temperatures (under weak regularization). Over 24 different problem sets, the combined MDOT-PNCG algorithm outperforms aggressively tuned Sinkhorn iteration and many other recent baselines in terms of convergence of the primal suboptimality gap measured in wall-clock time. The algorithm was also shown to behave well with respect to the problem size. Interesting directions for future research include the theoretical convergence behavior of PNCG, bounds on the gradient norm of our warm-started initialization, better warm-starting methods, and adaptive dual problem stopping criteria. The development of faster KL projection algorithms and adaptive temperature decay schedules are also of interest; see also our recent work in this direction (Kemertas et al., 2025).
Acknowledgements
@@
Amir-massoud Farahmand acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) through the Discovery Grant program (2021-03701). Mete Kemertas acknowledges the support of NSERC through the CGS-D scholarship. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Appendix
The Appendix is organized as follows:
  • Appendix A provides the proofs of our theoretical results.
  • Appendix B presents the line search routine we used for Alg. 2 after a primer on line search.
  • Appendix C investigates our use of $H_{\min}({\bm{r}},{\bm{c}})$ when deciding the KL projection stopping criteria in contrast to the more standard use of $\log n$.
  • Appendix D ablates the design decision in MDOT (Alg. 1) to vary the KL projection stopping criterion based on the present temperature, and shows its performance benefit.
  • Appendix E compares MDOT-PNCG to the network simplex solver in higher dimensions ($n=16384$) than the experiments in the main text ($n=4096$), and shows better scaling.
  • Appendix F discusses implementation details for baseline algorithms.
  • Appendix G provides experiments vs. the IPOT algorithm of Xie et al. (2020).
  • Appendix H provides additional benchmarking on 20 problem sets from the DOTmark benchmark (Schrieber et al., 2017).
Appendix A Proofs
Here, we provide proofs for the theoretical results in the main text; Lemma 4.1 in Appx. A.1, Proposition 4.2 in Appx. A.2, and Proposition 4.5 in Appx. A.3.
A.1 Proof of Lemma 4.1
See 4.1
Proof.
First, we emphasize the following observation, from which Lemma 4.1 follows immediately using ${P({\bm{u}},{\bm{v}};\gamma)\odot\exp\{-\Delta_{\gamma}C\}=P({\bm{u}},{\bm{v}};\gamma+\Delta_{\gamma})}$.
Observation A.1.
{mdframed}
[ skipabove=4pt, skipbelow=4pt, innertopmargin=3pt, innerbottommargin=3pt] Given any initial ${{\bm{u}},{\bm{v}}\in\mathbb{R}^{n}}$ and $P({\bm{u}},{\bm{v}};\gamma)$ defined as in (1), we have
$P^{*}(\gamma)=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}D_{\mathrm{KL}}\Big{(}P|P({\bm{u}},{\bm{v}};\gamma)\Big{)}=P({\bm{u}}^{*},{\bm{v}}^{*};\gamma),\textrm{ where }{\bm{u}}^{*},{\bm{v}}^{*}\in\operatorname*{arg\,min}g(\gamma,{\bm{r}},{\bm{c}}).$ (18)
In words, (i) minimizing the objective $g$ in (2.1) given any initial ${\bm{u}},{\bm{v}}\in\mathbb{R}^{n}$ amounts to a KL projection of $P({\bm{u}},{\bm{v}};\gamma)$ onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$, and (ii) the set of matrices $P({\bm{u}},{\bm{v}};\gamma)=\exp\{{\bm{u}}\bm{1}^{\top}+\bm{1}{\bm{v}}^{\top}-\gamma C\}$ all have the same KL projection in ${\mathcal{U}}({\bm{r}},{\bm{c}})$. The intuition here is that since the elements $u_{i},v_{j}$ of ${\bm{u}},{\bm{v}}$ simply rescale the $i^{\mathrm{th}}$ row and $j^{\mathrm{th}}$ column of $P({\bm{u}},{\bm{v}};\gamma)$, and the unique projection onto ${\mathcal{U}}({\bm{r}},{\bm{c}})$ is the optimal scaling (with ${\bm{r}}(P)={\bm{r}}$ and ${\bm{c}}(P)={\bm{c}}$), the specific initial scaling of rows and columns in the conditioning matrix $P({\bm{u}},{\bm{v}};\gamma)$ is irrelevant.
We now start the formal proof by deriving the relationship
$P({\bm{u}}^{*},{\bm{v}}^{*};\gamma)=\operatorname*{arg\,min}_{P\in{\mathcal{U}}({\bm{r}},{\bm{c}})}\left[D_{\mathrm{KL}}\big{(}P|P({\bm{u}},{\bm{v}};\gamma)\big{)}+K\right],$ (19)
given any initial ${\bm{u}},{\bm{v}}\in\mathbb{R}^{n}$ and ${\bm{u}}^{*},{\bm{v}}^{*}\in\operatorname*{arg\,min}_{{\bm{u}},{\bm{v}}\in\mathbb{R}^{n}}g({\bm{u}},{\bm{v}};\gamma)$. Here we have introduced a constant $K$ which, of course, does not effect the $\operatorname*{arg\,min}$. We will choose $K$ to simplify the derivation below. Note that $K$ can depend on the given quantities, such as ${\bm{u}}$, ${\bm{v}}$, and so on, but not on the unknown $P$. Given the definition of $D_{\mathrm{KL}}$ in Sec. 2 for un-normalized $P\in\mathbb{R}^{n\times n}_{>0}$, observe that
$D_{\mathrm{KL}}\big{(}P|P({\bm{u}},{\bm{v}};\gamma)\big{)}$ $=\langle P,\log P-\log P({\bm{u}},{\bm{v}};\gamma)\rangle+\langle P({\bm{u}},{\bm{v}};\gamma)-P,\bm{1}\rangle$
$=\langle P,\log P-{\bm{u}}\bm{1}^{\top}-\bm{1}{{\bm{v}}}^{\top}+\gamma C\rangle+\langle P({\bm{u}},{\bm{v}};\gamma)-P,\bm{1}\rangle$
$=\langle P,-\bm{1}+\gamma C+\log P\rangle-\langle{\bm{u}},{\bm{r}}(P)\rangle-\langle{\bm{v}},{\bm{c}}(P)\rangle+\underbrace{\left\lVert P({\bm{u}},{\bm{v}};\gamma)\right\rVert_{1}}_{\textrm{constant in }P}.$
As shown below, it is convenient to use $K=\langle{\bm{u}},{\bm{r}}\rangle+\langle{\bm{v}},{\bm{c}}\rangle-\left\lVert P({\bm{u}},{\bm{v}};\gamma)\right\rVert_{1}$, for which we find
$D_{\mathrm{KL}}\big{(}P|P({\bm{u}},{\bm{v}};\gamma)\big{)}+K=\langle P,-\bm{1}+\gamma C+\log P\rangle-\langle{\bm{u}},{\bm{r}}(P)-{\bm{r}}\rangle-\langle{\bm{v}},{\bm{c}}(P)-{\bm{c}}\rangle.$ (20)
We represent the equality constraints on $P$, that is, ${\bm{r}}(P)={\bm{r}}$ and ${\bm{c}}(P)={\bm{c}}$, using Lagrange multipliers $\bm{\alpha}$ and $\bm{\beta}$ to form the Lagrangian